Solve: 4x/3-3x+1/2=2 and 2 fifths
Answer by stanbon(48558)

Solve: 4x/3-3x+1/2=2 and 2 fifths
----------
(4/3)x - 3x + (1/2) = 12/5
(4/3)x - (9/3)x = (12/5)-(1/2)
(-5/3)x = (19/10)
x = (-3/5)(19/10)
x = -57/50
----------------
Cheers,
Stan H.